OpenStudy (anonymous):

Write an explicit definition for the sequence −2, 1, 4, 7,

jimthompson5910 (jim_thompson5910):

Hint: you're adding 3 to each term to get the next term ex: -2+3 = 1 1+3 = 4

OpenStudy (anonymous):

so would the answer be a^n = –2 + 3n

jimthompson5910 (jim_thompson5910):

close, but you're most likely starting off at n = 1

jimthompson5910 (jim_thompson5910):

so you have to subtract 3 to get -2 - 3 = -5 so if you're starting at n = 1, then the formula is a(n) = -5 + 3n or a(n) = 3n - 5

OpenStudy (anonymous):

hmm but the only choices i see are a(n) = –2 + 3n a(n) = –2 + (n – 3) a(n) = –2 + (n – 3) a(n) = –2 + 3 (n – 1)

jimthompson5910 (jim_thompson5910):

oh then you use this formula a(n) = a1 + d(n-1)

jimthompson5910 (jim_thompson5910):

a1 = -2 d = 3